OpenStudy (anonymous):

Solve the equation (2+sin A)/(2-cos A) = 2 where 0 <= A <= 180. Given the answers correct to the near 0.1 if necessary.

OpenStudy (welshfella):

lets this way cross multiply:- 2 + sin A = 4 - 2 cos A sin A + 2 cos A = 2

OpenStudy (welshfella):

its been a long time since I did these hold on a second while a look it up

OpenStudy (welshfella):

OK you need to write this in the form R(sin(A + B) where B is the complementary angle

OpenStudy (welshfella):

SInA + 2cosA = RsinAcosB + RsinBcosA equating coefficients of sinA and cosA sinA; 1 = RcosB cosA: 2 = RsinB

OpenStudy (welshfella):

so tan B = RsinB / RcosB = 2 making B = 63.43 degrees and R = square root of the sum of the coefficients of cos A and SinA = sqrt(1^2+2^2) = sqrt5

OpenStudy (welshfella):

so we get sqrt5(A + 63.43) = 2

OpenStudy (welshfella):

from this you can find the value of A

OpenStudy (welshfella):

oops sorry that should be sqrt5 sin( A + 63.43) = 2

OpenStudy (welshfella):

lol - I'm looking at the original equation and I see you can get one solution just by inspection. sin 0 = 0 and cos 0 = 1 so plugging this in (2 + 0) / (2 - 1) = 2 so A = 0 is one solution

OpenStudy (welshfella):

thats the only solution 180 wont fit because cos 180 = -1